OpenStudy (jenniferjuice):

Which of the following is a solution of x2 + 6x = −18? x = 3 − 3i x = −3 + 3i x = −6 + 3i x = 6 − 3i

OpenStudy (jenniferjuice):

@ramit.dour

OpenStudy (anonymous):

x=−3+3i,−3−3i

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x^2 + 6x = −18 }\) \(\normalsize\color{blue}{ x^2 + 6x +9= −9 }\) \(\normalsize\color{blue}{ (x+3)^2= −9 }\) \(\normalsize\color{blue}{ x+3= ±3i }\) \(\normalsize\color{blue}{ x= -3±3i }\)

OpenStudy (anonymous):

-3-3i

OpenStudy (solomonzelman):

and ` -3+3i `

OpenStudy (anonymous):

oh my bad

OpenStudy (solomonzelman):

well, you already said the correct answer, so no worries :)

OpenStudy (anonymous):

haha

OpenStudy (jenniferjuice):

:) thanks guys